# Discovering Preference Optimization Algorithms with and for Large Language Models

### v1

@@ A.1 Prompts
Upon a successful completion, we return the following user prompt, where ‘val’ is the MT-Bench score:
@@ B.1 Discovery Task - Single-turn Dialogue
For each valid generated objective function $f_{i}$, we use it to train an LLM and then collect a performance evaluation score. Specifically, we follow the same process when training and evaluating all objective functions, starting with a pre-trained supervised fine-tuned (SFT) 7 billion gemma model of ‘zephyr-7b-gemma-sft’ This is a 7 billion base version gemma [Gemma-Team et al., 2024] model supervised-fine-tuned on the ‘deita-10k-v0-sft’ dataset [Liu et al., 2023]. Starting with this model, we train it on the pairwise preference dataset of ‘Argilla DPO Mix 7K’; which attempts to create a high-quality preference dataset by filtering only highly rated chosen responses from the datasets of a multi-turn dataset, instruction following dataset [Longpre et al., 2023] and a diverse preference dataset that covers truthfulness, honesty and helpfulness [Cui et al., 2023]. For each training run, we trained all the parameters of the starting model, using a fixed $\beta=0.05$. We used the same fixed hyper-parameters for all training runs unless explicitly noted. Specifically, we used a learning rate of 5e-7, bfloat16 floating-point format, two epochs, a batch size per device of two, a gradient accumulation step of 8, a cosine learning rate scheduler, and AdamW optimization algorithm [Loshchilov and Hutter, 2017]. We use the popular TRL transformers library [vonwerra2022trl], adapting the offline preference optimization objective function to train all models. The models were trained on 8 nvidia A100 GPUs. An individual training run takes approximately 30 minutes. We provide training and evaluation statistics for discovered objective functions in Figure 7.
@@ B.2 TL;DR Summarization
To determine if the discovered objective functions generalize well also to other tasks, we use them to preference optimize an LLM for text summarization. Specifically, we start again with a pre-trained supervised fine-tuned (SFT) 7 billion gemma model of ‘zephyr-7b-gemma-sft’, and we optimized it with the objective function $f_{i}$ on a subsample of the the Reddit TL;DR summarization preference dataset [Völske et al., 2017] [^6] https://huggingface.co/datasets/CarperAI/openai_summarize_comparisons. More precisely we use the first 10% of the dataset for preference optimization, which ammounts to around 8’000 training samples. During training the hyperparameters are kept same as in the single turn dialogue task, explained in subsection B.1, except that LLMs where trained 4 nvidia A100 GPUS using a gradient accumulation step of 16. An individual training run takes approximately 1.5h.
@@ B.3 IMDb Positive Text Generation
Another popular generalization task for preference optimization [Rafailov et al., 2023] is to fine tune a small LLM to generate positive text for movie review, based on the IMDb sentiment dataset [Maas et al., 2011] [^7] https://huggingface.co/datasets/ZHZisZZ/imdb_preference. As starting model, we use a GPT2 model [Radford et al., 2019], that was supervised fine-tuned on the IMDb dataset [^8] https://huggingface.co/lvwerra/gpt2-imdb. Subsequently, we apply the baseline and discovered objective function $f_{i}$ for preference optimization. The goal of the LLM is given a short prompt of 2-8 tokens, which indicate the start of a movie review, to generate a positive review. As we are interested in the effect of $\beta$ on the rewards and KL-Divergence, we train the objective functions over a sweep of $\beta\in\{0.01,0.025,0.05,0.1,0.25,0.5,1,2.5,5\}$. Every the LLM is trained for three epochs, using AdamW optimizer, with an initial learning rate of 5.0e-5, a warm up scheduler of 0.1, a cosine learning rate scheduler. The models are trained on 4 nvidia A100 GPUs, using a gradient accumulation step of 8, and a batch size per device of 2. The training takes around 30 minutes.
@@ C.1 MT-Bench
To assess the fitness of the discovered preference optimization loss function during the discovery phase, we evaluate the trained LLMs on the MT-Bench [Zheng et al., 2024] benchmark. The evaluation benchmark consists of high quality 80 multi-turn questions, from various disciplines. The goal is to assess LLMs ability to follow instructions and keep the flow of a conversation. A larger LLM, in our case GPT-4, is then used as judge to score the quality of the answers with a number from 0 (lowest) to 10 (highest). Scores are given based on the quality of the LLMs first turn answer (single-turn), as well as on first and second answers (multi-turn). Finally, the MT-Bench score is the average of single-turn and multi-turn scores. For answer generation and evaluation, we used the FastChat library [^9] https://github.com/lm-sys/FastChat and its standard sampling and temperature parameters, provided by Zheng et al. [2024].
@@ C.2 Alpaca Eval
Currently, Alpaca Eval 2.0 [Li et al., 2023, Dubois et al., 2023, 2024] is also a popular benchmark to evaluate LLMs. This is a single-turn dialogue LLM-based automatic evaluation using a stronger LLM, here GPT-4 Turbo, to assess the win rate of the trained LLM policy’s completion compared either GPT-4 or to the of the underlying SFT base model. Specifically, Alpaca Eval 2.0, has been validated against 20K human annotations, and aims to reduce the length bias of Alpaca Eval; where using length controlled (LC) Alpaca Eval shows a correlation with Chatbot Arena of 0.98, making it a popular benchmark with the highest correlation to Chatbot Arena [Dubois et al., 2024]. The Alpaca evaluation dataset consists of 841 high quality instruction, originating from different data sets. The library [^10] https://github.com/tatsu-lab/alpaca_eval provided by Dubois et al. [2024] calculates the win-rate (percentage were the trained policy is prefered over the reference policy, first introduced in Alpaca Eval 1.0), and a length-controlled win-rate, where a linear model is fitted in order to de-bias for length of the prompt and instruction difficulty. To generate the answers we use a temperature of 0.7, sampling, and a maximum number of new tokens of 1024. Furthermore, the library provides the standard error of the mean, which indicates the confidence of the win-rate and LC win-rate.
@@ C.4 IMDb Rewards vs KL-Divergence
For the positive text generation, we do not require an LLM judge compared to MT-Bench, Alpaca Eval 2.0, and TL;DR evaluation, as we take a pre-trained sentiment classifier [^11] https://huggingface.co/siebert/sentiment-roberta-large-english as ground truth reward scorer. For the positive text generation, the LLMs apply sampling, and a maximum of 60 new tokens. The rewards and KL-divergence are averaged over 10 different generations from the trained LLMs.
@@ D.1 Frontiers of Expected Reward vs KL Divergence
D.2 Loss Sweeps for Different $\beta$ Parameters
@@ D.4 Visual Language Models for Objective Discovery
Appendix E Discovered Objective $f$ Functions
@@
Moreover, we display the code of the discovered losses as it is output by the LLM. In addition, we provide a mathematical representation of each, which we have adapted to be consistent with $\beta$ being the KL-Divergence regularization parameter. This is due to the fact, that the generated code for LRML, DBAQL, AQL, AQFL, and PFL did not uphold the $\beta$ ought to be multiplied with the difference of log-ratios, before any further calculations. If this was not uphold, it could to the loss function changing shapes based on the KL-regularization term, and therefore models could not converge, or potentially collapse. In future work, we should constrain the exploring LLM to uphold the $\beta$ multiplication with the input, before any other calculations are done with the difference of log-ratios $\rho$. As the meta exploration was done with a set $\beta=0.05$, and we wish to keep consistent with this scale of regularization, we have adapted the losses by dividing $\rho$ values used in intermediate calculations with a scalar $\tau=0.05$.
@@ E.1 DBAQL: Dynamic Blended Adaptive Quantile Loss
$f_{dbaql}(\beta\rho)$ $=\sigma(\mathds{V}\text{ar}[\rho/\tau])\cdot f_{dpo}(\beta\rho/0.9)+(1-\sigma(\mathds{V}\text{ar}[\rho/\tau]))\cdot f_{exp}(\beta\rho\cdot 0.9)$ (9)
@@ E.2 AQL: Adaptive Quantile Loss
$q$ $=\sigma(-\beta\cdot(\rho/\tau-m_{2}))$ (12)
$m_{2}$ $=0.5+0.01\cdot\big{(}\mathds{E}[\sigma(\rho/\tau)]-0.5\big{)}$ (13)
@@ E.4 AQFL: Adaptive Quantile Feedback Loss
$d$ $=|\rho\tau-m_{2}|$ (22)
$m_{2}$ $=m_{1}+0.05\cdot\big{(}\sigma(\mathds{E}[\rho/\tau]-m_{1})\big{)}$ (23)
$m_{1}$ $=\mathds{E}[\sigma(-\rho/\tau)]\cdot\sqrt{\mathds{V}\text{ar}[\rho/\tau]}$ (24)
@@ E.6 LRML: Log Ratio Modulated Loss
$f_{lrml}(\beta\rho)$ $=(1-\sigma(\rho/\tau))\cdot f_{dpo}(\beta\rho)+\sigma(\rho/\tau)\cdot f_{exp}(-\beta\rho)$ (27)
@@ E.7 PFL: Policy Focused Loss
Interestingly, the PFL generated function code did not include any beta values in the loss function. We have added it to the corrected code for the IMDb experiment, as well as to the mathematical expression below
$f_{pfl}(\beta\rho)$ $=1/2\cdot f_{dpo}(\beta\rho)\cdot\mathds{1}_{[\pi_{w}>\pi_{r}]}+2\cdot f_{slic}(\beta\rho)\cdot\mathds{1}_{[\pi_{w}\leq\pi_{r}]}$ (29)

### v2

@@ A.1 Prompts
Upon successful completion, we return the following user prompt, where ‘val’ is the MT-Bench score:
@@ B.1 Discovery Task - Single-turn Dialogue
For each valid generated objective function $f_{i}$, we use it to train an LLM and then collect a performance evaluation score. Specifically, we follow the same process when training and evaluating all objective functions, starting with a pre-trained supervised fine-tuned (SFT) 7 billion gemma model of ‘zephyr-7b-gemma-sft’ This is a 7 billion base version gemma [Gemma-Team et al., 2024] model supervised-fine-tuned on the ‘deita-10k-v0-sft’ dataset [Liu et al., 2023]. Starting with this model, we train it on the pairwise preference dataset of ‘Argilla DPO Mix 7K’; which attempts to create a high-quality preference dataset by filtering only highly rated chosen responses from the datasets of a multi-turn dataset, instruction following dataset [Longpre et al., 2023] and a diverse preference dataset that covers truthfulness, honesty and helpfulness [Cui et al., 2023]. For each training run, we trained all the parameters of the starting model, using a fixed $\beta=0.05$. We used the same fixed hyper-parameters for all training runs unless explicitly noted. Specifically, we used a learning rate of 5e-7, bfloat16 floating-point format, two epochs, a batch size per device of two, a gradient accumulation step of 8, a cosine learning rate scheduler, and AdamW optimization algorithm [Loshchilov and Hutter, 2017]. We use the popular TRL transformers library [von Werra et al., ], adapting the offline preference optimization objective function to train all models. The models were trained on 8 Nvidia A100 GPUs. An individual training run takes approximately 30 minutes. We provide training and evaluation statistics for discovered objective functions in Figure 7.
@@ B.2 TL;DR Summarization
To determine if the discovered objective functions generalize well also to other tasks, we use them to preference optimize an LLM for text summarization. Specifically, we start again with a pre-trained supervised fine-tuned (SFT) 7 billion gemma model of ‘zephyr-7b-gemma-sft’, and we optimized it with the objective function $f_{i}$ on a subsample of the Reddit TL;DR summarization preference dataset [Völske et al., 2017] [^6] https://huggingface.co/datasets/CarperAI/openai_summarize_comparisons. More precisely we use the first 10% of the dataset for preference optimization, which amounts to around 8’000 training samples. During training the hyperparameters are kept the same as in the single-turn dialogue task, explained in subsection B.1, except that LLMs were trained 4 Nvidia A100 GPUS using a gradient accumulation step of 16. An individual training run takes approximately 1.5 hours.
@@ B.3 IMDb Positive Text Generation
Another popular generalization task for preference optimization [Rafailov et al., 2023] is to fine-tune a small LLM to generate positive text for movie review, based on the IMDb sentiment dataset [Maas et al., 2011] [^7] https://huggingface.co/datasets/ZHZisZZ/imdb_preference following the the work of Zhou et al. [2024]. As a starting model, we use a GPT2 model [Radford et al., 2019], that was supervised fine-tuned on the IMDb dataset [^8] https://huggingface.co/lvwerra/gpt2-imdb. Subsequently, we apply the baseline and discovered objective function $f_{i}$ for preference optimization. The goal of the LLM is to give a short prompt of 2-8 tokens, which indicate the start of a movie review, to generate a positive review. As we are interested in the effect of $\beta$ on the rewards and KL-Divergence, we train the objective functions over a sweep of $\beta\in\{0.01,0.025,0.05,0.1,0.25,0.5,1,2.5,5\}$. Every LLM is trained for three epochs, using the AdamW optimizer, with an initial learning rate of 5.0e-5, a warm-up scheduler of 0.1, and a cosine learning rate scheduler. The models are trained on 4 Nvidia A100 GPUs, using a gradient accumulation step of 8, and a batch size per device of 2. The training takes around 30 minutes.
@@ C.1 MT-Bench
To assess the fitness of the discovered preference optimization loss function during the discovery phase, we evaluate the trained LLMs on the MT-Bench [Zheng et al., 2024] benchmark. The evaluation benchmark consists of high quality 80 multi-turn questions, from various disciplines. The goal is to assess LLM’s ability to follow instructions and keep the flow of a conversation. A larger LLM, in our case GPT-4, is then used as judge to score the quality of the answers with a number from 0 (lowest) to 10 (highest). Scores are given based on the quality of the LLM’s first-turn answer (single-turn), as well as on first and second answers (multi-turn). Finally, the MT-Bench score is the average of single-turn and multi-turn scores. For answer generation and evaluation, we used the FastChat library [^9] https://github.com/lm-sys/FastChat and its standard sampling and temperature parameters, provided by Zheng et al. [2024].
@@ C.2 Alpaca Eval
Currently, Alpaca Eval 2.0 [Li et al., 2023, Dubois et al., 2023, 2024] is also a popular benchmark to evaluate LLMs. This is a single-turn dialogue LLM-based automatic evaluation using a stronger LLM, here GPT-4 Turbo, to assess the win rate of the trained LLM policy’s completion compared to either GPT-4 or the of the underlying SFT base model. Specifically, Alpaca Eval 2.0, has been validated against 20K human annotations, and aims to reduce the length bias of Alpaca Eval; where using length-controlled (LC) Alpaca Eval shows a correlation with Chatbot Arena of 0.98, making it a popular benchmark with the highest correlation to Chatbot Arena [Dubois et al., 2024]. The Alpaca evaluation dataset consists of 841 high-quality instructions, originating from different data sets. The library [^10] https://github.com/tatsu-lab/alpaca_eval provided by Dubois et al. [2024] calculates the win-rate (percentage were the trained policy is prefered over the reference policy, first introduced in Alpaca Eval 1.0), and a length-controlled win-rate, where a linear model is fitted to de-bias for length of the prompt and instruction difficulty. To generate the answers we use a temperature of 0.7, sampling, and a maximum number of new tokens of 1024. Furthermore, the library provides the standard error of the mean, which indicates the confidence of the win-rate and LC win-rate.
@@ C.4 IMDb Rewards vs KL-Divergence
For the positive text generation, we do not require an LLM judge compared to MT-Bench, Alpaca Eval 2.0, and TL;DR evaluation, as we take a pre-trained sentiment classifier [^11] https://huggingface.co/siebert/sentiment-roberta-large-english as ground truth reward scorer. For the positive text generation, the LLMs apply sampling and a maximum of 60 new tokens. The rewards and KL-divergence are averaged over 10 different generations from the trained LLMs.
@@ D.1 Frontiers of Expected Reward vs KL Divergence
D.2 Loss Sweeps for Different Beta Parameters
@@ D.4 Visual Language Models for Objective Discovery
Appendix E Discovered Objective Functions
@@
Moreover, we display the code of the discovered losses as it is output by the LLM. In addition, we provide a mathematical representation of each, which we have adapted to be consistent with $\beta$ being the KL-Divergence regularization parameter. This is due to the fact, that the generated code for LRML, DBAQL, AQL, AQFL, and PFL did not uphold the $\beta$ ought to be multiplied with the difference of log-ratios, before any further calculations. If this was not upheld, it could to the loss function changing shapes based on the KL-regularization term, and therefore models could not converge, or potentially collapse. In future work, we should constrain the exploring LLM to uphold the $\beta$ multiplication with the input, before any other calculations are done with the difference of log-ratios $\rho$. As the meta exploration was done with a set $\beta=0.05$, and we wish to keep consistent with this scale of regularization, we have adapted the losses by dividing $\rho$ values used in intermediate calculations with a scalar $\tau=0.05$.
@@ E.1 DBAQL: Dynamic Blended Adaptive Quantile Loss
$f_{dbaql}(\beta\rho)$ $=\sigma(\mathds{V}\text{ar}[\beta\rho/\tau])\cdot f_{dpo}(\beta\rho/0.9)+(1-\sigma(\mathds{V}\text{ar}[\beta\rho/\tau]))\cdot f_{exp}(\beta\rho\cdot 0.9)$ (9)
@@ E.2 AQL: Adaptive Quantile Loss
$q$ $=\sigma(\tau m_{2}-\beta\rho)$ (12)
$m_{2}$ $=0.5+0.01\cdot\big{(}\mathds{E}[\sigma(\beta\rho/\tau)]-0.5\big{)}$ (13)
@@ E.4 AQFL: Adaptive Quantile Feedback Loss
$d$ $=|\beta\rho/\tau-m_{2}|$ (22)
$m_{2}$ $=m_{1}+0.05\cdot\big{(}\sigma(\mathds{E}[\beta\rho/\tau]-m_{1})\big{)}$ (23)
$m_{1}$ $=\mathds{E}[\sigma(-\beta\rho/\tau)]\cdot\sqrt{\mathds{V}\text{ar}[\beta\rho/\tau]}$ (24)
@@ E.6 LRML: Log Ratio Modulated Loss
$f_{lrml}(\beta\rho)$ $=(1-\sigma(\beta\rho/\tau))\cdot f_{dpo}(\beta\rho)+\sigma(\beta\rho/\tau)\cdot f_{exp}(-\beta\rho)$ (27)
@@ E.7 PFL: Policy Focused Loss
Interestingly, the PFL generated function code did not include any $\beta$ values in the loss function. We have added it to the corrected code for the IMDb experiment, as well as to the mathematical expression below. We account for the consistency of the logit values by dividing with $\tau=0.05$, which is the same value as $\beta$ in the discovery process.
$f_{pfl}(\beta\rho)$ $=1/2\cdot f_{dpo}(\beta\rho/\tau)\cdot\mathds{1}_{[\pi_{w}>\pi_{r}]}+2\cdot f_{slic}(\beta\rho/\tau)\cdot\mathds{1}_{[\pi_{w}\leq\pi_{r}]}$ (29)
$\tau=0.05$ (30)